OpenStudy (anonymous):

what is the exact value of the expression sqrt486- sqrt +sqrt6? simplify if possible 2sqrt2 2sqrt6 6sqrt2 8sqrt6

OpenStudy (anonymous):

Could you please write the problem again? "sqrt486- sqrt +sqrt6" There is an sqrt with nothing inside.

OpenStudy (anonymous):

\[\sqrt{486}-\sqrt{24}+\sqrt{6}\]

OpenStudy (anonymous):

24=4*6 So then: sqrt(24)=sqrt(4)*sqrt(6)=2*sqrt(6) What can we do with the 486?

OpenStudy (anonymous):

54=9*6 I think that you can go on by your own. Give it a try!

OpenStudy (anonymous):

\[\sqrt{54*9}-\sqrt{6*4}+\sqrt{6}\] \[3\sqrt{54}-2\sqrt6+\sqrt6\]

OpenStudy (anonymous):

\[6\sqrt{2}\]

OpenStudy (anonymous):

\[3\sqrt{54}-2\sqrt{6}+\sqrt{6}=9\sqrt{6}-2\sqrt{6}+\sqrt{6}\] Agree with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So.. What do we get?

OpenStudy (anonymous):

\[8\sqrt{6} thank you for your help!\]

OpenStudy (anonymous):

You are welcome! Glad to help! :)